How do I find the answer to: Joanne invested four times as much money in an account paying 5% as she did in an account paying 6%. The total amount of interest for the year was $520. How much money did Joanne invest in each account?
Answer by mananth(16946)

let amount invested at 6% be x
amount invested at 5% will be 4x

Interest = 520
6%x+5%*4x = 520
multiply by 100
6x+20x= 52000
26x=52000
/26
x= 2000 @ 6%
4*2000 = 8000 at 5%